OpenStudy (anonymous):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 3x - 2.

zepdrix (zepdrix):

\[\large\rm domain=\{\color{orangered}{-2,0,2}\}\] Hey there :) So our domain values go `into the function`. Here is the first one:\[\large\rm y=3(\color{orangered}{-2})-2\]\[\large\rm y=-6-2\]\[\large\rm y=-8\]

zepdrix (zepdrix):

This `output value` is part of our range. So we have the first value for our range,\[\large\rm range=\{\color{orangered}{-8,~~~,~~~}\}\]

zepdrix (zepdrix):

Think you can try that with the other values from the domain? :)